# Measuring What Matters!! Assessing Therapeutic Principles in Mental-Health Conversation

### 07 Apr 2026 v1

@@ Appendix D Experimental Setup
Model Acc. Macro Weighted
F1 Prec. Rec. F1 Prec. Rec.
Dataset: PTSD
GPT-4o 38.26 25.53 28.79 28.49 35.06 37.09 38.26
Qwen 38.00 16.80 19.10 25.30 31.10 34.90 38.00
Llama 3.2 29.80 8.80 8.60 18.90 17.50 14.50 29.80
CARE 52.34 46.80 48.80 46.50 50.90 51.30 52.40
Dataset: CheeseBurger
GPT-4o 26.33 12.62 19.61 10.14 31.50 42.22 26.33
Qwen 28.05 11.25 10.71 16.58 26.94 29.35 28.05
Llama 3.2 23.15 9.01 16.98 17.89 18.95 39.13 23.15
CARE 48.96 20.46 28.28 20.07 50.30 65.25 48.96
Instruction: I need you to think like a mental health counselor and use a hidden chain-of-thought process. I will provide you with a therapist dialogue along with several reference dialogues. First, internally analyze and understand the reference dialogues to build expert knowledge. Then, using your hidden chain-of-thought reasoning, analyze the therapist dialogue and develop a detailed explanation for Active Listening. Finally, provide your explanation in exactly 100 words. Do not include any of your internal reasoning in the final output. Label Exclusive Utterances: It’s really important for us to consider that kind of thinking, hmm. Okay, Marsha, it’s good that we’ve got those things on the board. And what I’d like to do now is focus, stepping out a bit from your situation and looking, looking at it from another point of view. To me, it sounds like you’ve, you’re doing a whole lot of things really well. Can you list all of the things that you’re doing really well? Therapist Current Utterance: Okay, you’re doing great, just keep going.
Full Few-Shot In-Context Example Used for GPT-4o Evaluation Instruction: You are an expert in mental health counselling. Evaluate each therapist’s response on a scale from -2 to +2 based on the following six therapeutic principles: Non-Judgmental Language; Warmth and Encouragement; Respect for Autonomy; Active Listening; Reflecting Feelings; Situational Appropriateness. Use the reference examples below to guide your judgment. Each example includes a therapist response, detailed explanations for each principle, and final scores. Follow the same reasoning pattern when evaluating the new response. Example 51 Therapist Response: “So you have difficulty falling asleep. But once you fall asleep.” Explanations: Non-Judgmental Language: Non-judgmental language in therapy is essential for creating a safe and supportive environment where clients feel comfortable sharing their experiences without fear of criticism. Warmth and Encouragement: Warmth and encouragement in therapy create a safe space for clients to explore their feelings and challenges. Respect for Autonomy: Respect for autonomy emphasizes recognizing the client’s ability to make their own decisions regarding their mental health. Active Listening: Active listening involves fully concentrating on the client’s words and demonstrating understanding. Reflecting Feelings: Reflecting feelings acknowledges the client’s emotional experience by summarizing their struggle with sleep difficulties. Situational Appropriateness: The therapist response appropriately addresses a common clinical concern raised by the client. Scores: Non-Judgmental: 2, Warmth: 1, Respect: 1, Active: 2, Reflecting: 1, Situational: 2 Example 617 Therapist Response: “What County are we in?” Explanations: Non-Judgmental Language: The response does not express judgment and maintains neutrality. Warmth and Encouragement: The question provides limited warmth and encouragement. Respect for Autonomy: The response does not restrict the client’s autonomy. Active Listening: The response demonstrates minimal engagement with the client’s expressed concerns. Reflecting Feelings: The question serves as a grounding technique rather than emotional reflection. Situational Appropriateness: The response is weakly appropriate as it diverts attention from the client’s emotional needs. Scores: Non-Judgmental: 1, Warmth: 0, Respect: 1, Active: 1, Reflecting: 1, Situational: 1 Now evaluate this response: Therapist Response: “Okay, good. I’$l$ see you next week. And we’$l$ see how you did with those goals.” Explanations: Non-Judgmental Language: Non-judgmental language fosters a supportive therapeutic environment. The phrase “we’$l$ see how you did with those goals” avoids explicit criticism but implies evaluation. Warmth and Encouragement: The response conveys mild encouragement through continuity and reassurance. Respect for Autonomy: The therapist acknowledges the client’s responsibility for goal progress without imposing directives. Active Listening: The response reflects acknowledgment of prior discussion about goals and future follow-up. Reflecting Feelings: Emotional reflection is limited, as the response focuses on task completion rather than affect. Situational Appropriateness: The response is appropriate as a session-closing statement that reinforces accountability and continuity.
Appendix E Prompting with Label-Guided Reasoning
# In-Context Dim Macro Weighted
F1 Prec. Rec. F1 Prec. Rec.
GPT-4o Zero Shot NJ 23.67 25.81 25.75 23.67 25.81 25.75
W&E 20.44 33.60 24.55 20.44 33.60 24.55
RA 30.99 31.32 34.73 30.99 31.32 34.73
AL 37.20 42.19 33.53 37.20 42.19 33.53
RF 38.94 45.01 37.42 38.94 45.01 37.42
SA 31.68 39.18 30.53 31.68 39.18 30.53
Avg. 30.49 36.18 31.08 30.49 36.18 31.08
Few-Shot NJ 13.01 19.14 25.74 13.01 19.14 25.74
W&E 39.22 44.65 44.91 39.22 44.65 44.91
RA 18.52 45.83 27.24 18.52 45.83 27.24
AL 24.40 38.45 31.13 24.40 38.45 31.13
RF 18.27 27.95 23.65 18.27 27.95 23.65
SA 25.31 33.95 37.42 25.31 33.95 37.42
Avg. 23.12 34.99 31.69 23.12 34.99 31.69
Appendix F Human-Level Rational Assessment
Rationale Examples
Appendix G Additional Ablation Details.
Appendix H Behavioral Differences Between CARE and GPT-4o

### 13 Apr 2026 v2

@@ Appendix B Experimental Setup
Appendix C Additional Ablation Details.
Few-Shot vs Zero-Shot Prompt Design for GPT-4o Evaluation Few-Shot Prompt: Instruction: You are a clinical NLP evaluator. Score a therapist response from -2 to +2 across six therapeutic principles. You will be shown few-shot examples and then asked to score a new interaction. ### Example Patient: <previous utterance> Therapist: <therapist response> Explanations: Non-Judgmental Language: ... Warmth and Encouragement: ... Respect for Autonomy: ... Active Listening: ... Reflecting Feelings: ... Situational Appropriateness: ... Scores: NJ: X, W: X, RA: X, AL: X, RF: X, SA: X Repeat for multiple examples. ### Now evaluate this response: Patient: <current patient utterance> Therapist: <current therapist response> Explanations: ... Return scores in same format. Zero-Shot Prompt: Instruction: You are a mental health evaluation assistant trained in psychotherapy communication techniques. Please assess the quality of the therapist’s response using the following six therapeutic dimensions. Use the preceding patient’s statement to help you understand the context and appropriateness of the therapist’s message. Patient said: "<patient utterance>" Therapist responded: "<therapist utterance>" Rate the therapist’s response from -2 to +2 for each dimension: - -2 = strongly violates the principle - 0 = neutral or unclear - +2 = strongly demonstrates the principle Output Format (verbatim): Non-Judgmental: X, Warmth: X, Respect: X, Active Listening: X, Reflecting Feelings: X, Situational Appropriateness: X
Prompt Design for Dimension Specific COT Explanation Instruction: I need you to think like a mental health counselor and use a hidden chain-of-thought process. I will provide you with a therapist dialogue along with several reference dialogues. First, internally analyze and understand the reference dialogues to build expert knowledge. Then, using your hidden chain-of-thought reasoning, analyze the therapist dialogue and develop a detailed explanation for Active Listening. Finally, provide your explanation in exactly 100 words. Do not include any of your internal reasoning in the final output. Label Exclusive Utterances: It’s really important for us to consider that kind of thinking, hmm. Okay, Marsha, it’s good that we’ve got those things on the board. And what I’d like to do now is focus, stepping out a bit from your situation and looking, looking at it from another point of view. To me, it sounds like you’ve, you’re doing a whole lot of things really well. Can you list all of the things that you’re doing really well? Therapist Current Utterance: Okay, you’re doing great, just keep going.
Full Few-Shot In-Context Example Used for GPT-4o Evaluation Instruction: You are an expert in mental health counselling. Evaluate each therapist’s response on a scale from -2 to +2 based on the following six therapeutic principles: Non-Judgmental Language; Warmth and Encouragement; Respect for Autonomy; Active Listening; Reflecting Feelings; Situational Appropriateness. Use the reference examples below to guide your judgment. Each example includes a therapist response, detailed explanations for each principle, and final scores. Follow the same reasoning pattern when evaluating the new response. Example A Therapist Response: “So you have difficulty falling asleep. But once you fall asleep.” Explanations: Non-Judgmental Language: Non-judgmental language in therapy is essential for creating a safe and supportive environment where clients feel comfortable sharing their experiences without fear of criticism. Warmth and Encouragement: Warmth and encouragement in therapy create a safe space for clients to explore their feelings and challenges. Respect for Autonomy: Respect for autonomy emphasizes recognizing the client’s ability to make their own decisions regarding their mental health. Active Listening: Active listening involves fully concentrating on the client’s words and demonstrating understanding. Reflecting Feelings: Reflecting feelings acknowledges the client’s emotional experience by summarizing their struggle with sleep difficulties. Situational Appropriateness: The therapist response appropriately addresses a common clinical concern raised by the client. Scores: Non-Judgmental: 2, Warmth: 1, Respect: 1, Active: 2, Reflecting: 1, Situational: 2 Example B Therapist Response: “What County are we in?” Explanations: Non-Judgmental Language: The response does not express judgment and maintains neutrality. Warmth and Encouragement: The question provides limited warmth and encouragement. Respect for Autonomy: The response does not restrict the client’s autonomy. Active Listening: The response demonstrates minimal engagement with the client’s expressed concerns. Reflecting Feelings: The question serves as a grounding technique rather than emotional reflection. Situational Appropriateness: The response is weakly appropriate as it diverts attention from the client’s emotional needs. Scores: Non-Judgmental: 1, Warmth: 0, Respect: 1, Active: 1, Reflecting: 1, Situational: 1 Now evaluate this response: Therapist Response: “Okay, good. I’$l$ see you next week. And we’$l$ see how you did with those goals.” Explanations: Non-Judgmental Language: Non-judgmental language fosters a supportive therapeutic environment. The phrase “we’$l$ see how you did with those goals” avoids explicit criticism but implies evaluation. Warmth and Encouragement: The response conveys mild encouragement through continuity and reassurance. Respect for Autonomy: The therapist acknowledges the client’s responsibility for goal progress without imposing directives. Active Listening: The response reflects acknowledgment of prior discussion about goals and future follow-up. Reflecting Feelings: Emotional reflection is limited, as the response focuses on task completion rather than affect. Situational Appropriateness: The response is appropriate as a session-closing statement that reinforces accountability and continuity.
Appendix D Prompting
D.1 Few-shot vs Zero-shot
The few-shot setting provides multiple in-context examples with detailed explanations, enabling the model to learn implicit scoring patterns and align its judgments accordingly. In contrast, the zero-shot prompt relies solely on task instructions and the immediate dialogue context, without prior demonstrations. This distinction highlights how contextual grounding through examples can influence model reasoning, consistency, and interpretability in therapeutic assessment tasks. Overall, Figure 6 emphasizes the trade-off between guided learning in few-shot prompting and the generalization capability of zero-shot evaluation.
D.2 Prompting with Label-Guided Reasoning
Appendix E Human-Level Rational Assessment
Rationale Examples
Appendix F Behavioral Differences Between CARE and GPT-4o